1. A cellular phone company charges $60 per month for a service plan that includes 500 free text messages. In addition to the basic monthly rate, the company charges $0.15 for each additional text message over 500. A customer on this plan receives a bill for $111.30. How many text messages did this customer send and receive during the month?

2. You have a choice of receiving a wage of $34,000 per year, $2940 per month, $690 per week, or $16.10 per hour. Which pay choice would you take? Assume a 40-hour work week and 52 weeks of work per year.
A. $34,000 per year
B. $2940 per month
C. $690 per week
D. $16.10 per hour
1
Expert's answer
2020-09-28T20:13:01-0400

1. 111.3-60=51.3

51.3/0.15=342

342+500=842

Answer: 842 messages.

2. 2940*12=35280

690*52=35880

16.1*40*52=33488

Answer:690$ per week, because it's a most paying variant.